OpenStudy (anonymous):

3/y = y/5y-12 answer: -13, 3?? 3/y = y/5y-12 answer: -13, 3?? @Mathematics

OpenStudy (agreene):

\[\frac3y = \frac{y}{5y-12}\] is that the question?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i meant -12,3

OpenStudy (agreene):

I think it's positive 12 and 3. After completing the square, I have \[y-\frac{15}{12}=\pm\frac{9}{2}\] which should give y=3 and y = 12